# Seeing World Dynamics in a Nutshell

### v1

@@ 1 Introduction
Our natural world exhibits inherent dynamism —-from rustling leaves swaying in the wind to clouds drifting across the sky and ocean waves rolling along the shore—where objects maintain their structural integrity while undergoing continuous spatiotemporal evolution. A fundamental objective in video processing [65, 5] is to enable machines to interpret such visual information, allowing recovery of both object geometry and motion patterns while preserving the spatiotemporal coherence intrinsic to human perception. This capability is crucial for numerous applications, ranging from autonomous driving [20, 34] and robotics [24, 25, 43] to augmented reality and content creation [61, 31, 2, 85], where an accurate understanding of dynamic scenes directly impacts system performance and user experience.
@@
The current neural-based video representation [44, 58, 82] focuses predominantly on 2D and 2.5D techniques, treating videos as collections of spatiotemporal pixels [17, 91]. Although these discrete representations enable basic temporal modeling through pixel matching [60] and tracking [27, 81], they struggle to capture complex scene dynamics and maintain geometric consistency, particularly in challenging scenarios involving occlusions and non-rigid deformations [42, 44]. Moreover, these approaches inherently lack explicit 3D understanding, leading to unreliable spatial arrangements and limitations in novel view synthesis [40].
@@
Drawing inspiration from monocular video as a projection of the dynamic 3D world, we investigate whether videos can be represented in their intrinsic 3D form without per-scene optimization. Recent advances in dynamic Gaussian Splatting [29] have shown unprecedented capabilities in dynamic scene reconstruction [80, 62, 75, 16], achieving high-fidelity rendering with explicit 3D representation and superior efficiency. By modeling videos as continuous flows of Gaussian primitives over time, we address the limitations of 2D representations while enabling efficient video processing. As shown in Figure 1, this paradigm treats space-time as an entirety and offers key advantages: each Gaussian functions as a flexible building block adapting to local geometric structures, enabling seamless representation of complex scenes across scales. Moreover, when embraced with dynamic attributes, these structured Gaussians naturally approximate the underlying spatiotemporal 4D volume of dynamic scenes, capturing intrinsic motions with demonstrated temporal consistency and facilitating various downstream tasks including novel view synthesis [40, 86], video editing [7, 49], and frame interpolation [14, 39].
@@
However, transforming casually captured monocular videos to dynamic Gaussian representations instantly (sec. / frame) presents several challenges: (1) Unposed inputs. Gaussian Splatting [29] and its variants [87, 88, 23] heavily rely on accurate camera poses obtained through Structure-from-Motion (SfM) [66, 54], which are typically unavailable for casual monocular videos. Without such pose guidance, current methods [19, 4, 35, 73] fail to disentangle camera motion from scene dynamics, resulting in deteriorated rendering quality and even collapsed reconstruction. (2) Nonstructural Nature. Most existing dynamic Gaussian Splatting either leverage per-scene optimized deformation networks [80, 75, 16, 28, 32, 79] or adopt per-frame Gaussian generation [53] for dynamics modeling, both incompatible with our feed-forward prediction paradigm. The spatially unstructured property of Gaussian primitives further makes them prone to local minima in inverse rendering [95, 10], impeding feed-forward modeling of dynamic 3D scenes. (3) Spatial Ambiguity. The absence of multi-view supervision and initialization from SfM points significantly limits the spatial modeling capability of Gaussian Splatting, leading to ambiguous scale, depth collapse, and inconsistent spatial arrangements in reconstructed scenes.
@@
To address the above challenges, we introduce NutWorld to efficiently transform monocular videos into dynamic Gaussian Splatting representations in a single forward pass. Our approach comprises three key components: (1) A structured spatial-temporal aligned Gaussian (STAG) representation in canonical space (Section 4.1), which enables feed-forward prediction while ensuring pose-free and scale-invariant modeling. (2) An optimization-free pipeline (Section 4.2) that learns spatial-temporal correspondence and continuous temporal dynamics across video frames, efficiently converting them into STAG representations. (3) Depth and flow regularization (Section 4.3) that incorporates calibrated monocular depth [78] and optical flow priors [76] to resolve spatial ambiguity and motion-appearance entanglement in the ill-posed monocular video setting [62, 69]. Through large-scale pre-training, NutWorld could process arbitrarily long videos while maintaining spatial-temporal consistency via segment-based inference (Section 4.4).
@@
We conducted qualitative and quantitative experiments on RealEstate10K [94] and MiraData [26] to verify the representation efficacy of our proposed NutWorld in terms of video reconstruction. Moreover, our approach demonstrates compelling flexibility while maintaining real-time inference speed across various video downstream applications, including novel view synthesis, consistent depth estimation, video segmentation, as well as video editing and frame interpolation, suggesting its potential as a general-purpose video representation framework. The contributions and novelties of our paper are summarized as follows:
@@
  • We present the first framework to efficiently represent world dynamics in casually captured monocular videos as dynamic Gaussian Splatting in a single forward pass.
@@
  • Extensive experiments on video reconstruction and various downstream tasks demonstrate the spatial-temporal coherence and the versatility of NutWorld.
@@ 2 Related Work
Dynamic 3D scene reconstruction. Dynamic scenes, with moving objects and varying lighting conditions, require models that can handle temporal variations and motion in 3D space. While early works extended NeRF [40] to dynamic scenes [48, 46, 47] through temporal representations [48] and deformable fields [6], they remain computationally intensive due to MLPs and volume rendering [41]. Recent dynamic 3D Gaussian Splatting addresses this limitation by using deformable 3D Gaussian [80, 75, 28, 32] or anisotropic 4D Gaussian [16, 79], enabling efficient rendering through tile-based rasterization [29]. Among them, concurrent works [69, 62, 71, 33] have demonstrated dynamic scene reconstruction from unposed monocular videos. However, they rely on computationally intensive per-scene optimization, making them impractical for real-world applications. In contrast, our NutWorld focuses on neural video representation (NVR) without the availability of camera poses, specifically addressing the challenge of explicitly representing monocular videos rather than dynamic 3D scene reconstruction. Moreover, it employs a feed-forward approach that enables fast inference with high reconstruction quality and temporal consistency, distinguishing itself from tedious optimization-based 3D reconstruction paradigms.
Feed-Forward Gaussian Splatting. Recent advance in large-scale 3D scene datasets [36, 94, 37] has enabled feed-forward Gaussian approaches [8, 9, 90, 92, 74, 84, 55], which excel in efficiency and sparse view reconstruction. PixelSplat [8] and LatentSplat [74] employ the epipolar transformer [21, 72] to establish cross-view correspondences and predict 3D Gaussians from multi-view image features, while MVSplat [9] utilizes cost volumes to jointly predict depth and Gaussian parameters. Alternatively, GS-LRM [92] and Long-LRM [96] consider Gaussian Splatting reconstruction as a sequence-to-sequence translation task, employing transformer-based [67] or hybrid [12] architectures to regress Gaussian primitives. Although these methods address static 3D reconstruction, NutWorld is the first to represent dynamic scenes from unposed monocular videos.
@@ 4.1 Spatial-Temporal Aligned Gaussian
Canonical camera space. Given an unposed monocular video, we employ an orthographic camera coordinate system instead of an absolute 3D world coordinate system. This choice is motivated by two key challenges: the difficulty of obtaining coherent camera trajectories in dynamic scenes [66, 54, 45, 70, 68], and the inherent scale ambiguity in feed-forward 3D reconstruction models [8, 9, 90], where perspective projection couples object scale with its distance from camera. By adopting orthographic projection with a fixed pose along the $z$ axis, we eliminate the necessity for explicit camera pose estimation while modeling both camera and object motion without scale ambiguity in a unified canonical space. We provide a detailed implementation of the orthographic rasterization pipeline in the Appendix.
Structured Dynamic Gaussian. To overcome the unstructured nature in dynamic Gaussian Splatting and facilitate neural network integration, we introduce Spatial-Temporal Aligned Gaussian Splatting (STAG) in the canonical camera space. STAG constrains each dynamic Gaussian to a specific pixel location and timestamp, in contrast to the previous approach of predicting unconstrained Gaussians with deformable fields across orthogonal space-time. Formally, for an input frame $F_{k}$ with normalized timestamp $t_{k}\in[0,1]$, we compute a Gaussian feature map $\mathcal{E}^{k}\in\mathbb{R}^{U\times V\times T}$, where $U$ and $V$ represent spatial dimensions, and $T$ denotes the channel dimension. Each $T$-dimensional pixel is decoded into a 3D Gaussian with an associated deformation field $\{\mathcal{G}_{i},\mu_{i}(t)\}$** Note that based on our empirical observation, we simplify the deformation modeling by considering only the center position, such that the deformation field reduces to $\mathcal{D}_{i}(t)=\mu_{i}(t)$. in a pixel-aligned manner [8, 63].
@@
As shown in Figure 2, the temporal slicing function $\mathbbm{1}(t_{k},t_{j})$ modulates the deformation field $\mu^{k}(t_{j})$ based on the temporal relationship between the rendering timestamp $t_{j}$ and the Gaussian’s reference timestamp $t_{k}$. For $t_{j}=t_{k}$, the deformation is suppressed ($\mathbbm{1}(t_{k},t_{j})=0$), preserving the original Gaussian position $\mu^{k}$ to maintain spatial alignment. When $t_{j}\neq t_{k}$, the deformation field $\mu^{k}(t_{j})$ is activated to adjust $\mu^{k}$, allowing the Gaussian center to adapt to temporal differences.
@@ 4.2 Encapsulate Dynamics within “Nutshell”
This mapping generates a set of STAGs $\{\mathcal{G}_{i},\mu_{i}(t)\}$ that can be rendered into arbitrary frames ($M\geq K$) through temporal interpolation of the deformation field. Specifically,
Transformer-based Encoder. For each input frame $F_{k}\in\mathbb{R}^{H\times W\times 3}$, we augment pixels by concatenating their RGB values with corresponding depth coordinates $d^{*}_{k}$ along the channel dimension, obtained from a pre-trained depth estimation model [78]. The augmented frames are first split into non-overlapping patches of size $p\times p$ using convolutions, which are then linearly transformed and concatenated across all frames to generate transformer input tokens. Notably, our architecture eliminates the need for explicit positional embeddings as used in ViT [15, 1], since depth coordinates inherently encode spatial information. The transformer blocks, comprising self-attention [67] and MLP layers, process these concatenated tokens to capture spatiotemporal correspondence and produce encoded features $\mathcal{E}_{0}\in\mathbb{R}^{K\times h\times w\times C}$, where $h=H/p$ and $w=W/p$ denote the spatial resolution and $C$ denotes the feature dimension. To ensure sufficient STAGs for casual videos, we leverage a hierarchical upsampling network [92, 77] that progressively expands the spatial resolution of the encoded feature $\mathcal{E}_{0}$. Each upsampler block first expands the feature dimension by a factor of 4 through a linear layer, followed by a PixelShuffle [57] layer that doubles the spatial resolution. The resulting features are then processed by a local attention layer with window size $\mathcal{W}$, which balances computational efficiency and spatial-temporal feature aggregation:
@@ 4.3 Calibrated 2D Priors Regularization
Learning spatially-aware STAGs solely from monocular videos is inherently ill-posed due to depth ambiguity and motion uncertainty. Therefore, we leverage off-the-shelf foundation models [76, 78] to recover spatial relationships and temporal motion through calibrated depth and flow priors, respectively. Under the orthographic projection, the movement of STAG along the xy-coordinates directly corresponds to optical flow magnitude, whereas depth-related loss exclusively affects the z-coordinate, further facilitating the incorporation of those 2D priors.
Depth Regularization. To enhance robustness against scale and shift variations in depth rendering, we employ a scale and shift invariant loss [3, 52]. This loss function computes optimal scaling $\beta$ and shifting $\gamma$ factors that align the rendered depth $d$ with the pseudo depth $d^{*}$ estimated by the depth prior model [78]. The optimal values for $\beta$ and $\gamma$ are obtained by minimizing the squared error between the scaled rendered depth and the pseudo depth, as follows:
@@ 5.1 Experimental Setup
Training Dataset. NutWorld is pre-trained on MiraData [26] and RealEstate10K [94]. MiraData is a high-quality video dataset consisting primarily of 3D engine-generated scenes and movie clips with diverse motion patterns. The RealEstate10K dataset contains indoor house tour videos that showcase various architectural scenes and camera movements Unlike previous Generalizable 3DGS approaches [9, 8, 90], we treat RealEstate10K as a pure video dataset rather than a multi-view 3D scene dataset, thus not utilizing the COLMAP-calibrated camera poses. During pre-processing, we segment the original videos into video cubes, each containing 10 consecutive frames as the basic processing unit. Detailed information on the description of the dataset, preprocessing, and splitting of trains is provided in the Appendix.

### v2

@@ 1 Introduction
Our natural world exhibits inherent dynamism —-from rustling leaves swaying in the wind to clouds drifting across the sky and ocean waves rolling along the shore—where objects maintain their structural integrity while undergoing continuous spatiotemporal evolution. A fundamental objective in video processing [63, 5] is to enable machines to interpret such dynamic information, allowing recovery of both object geometry and motion patterns while preserving the spatiotemporal coherence intrinsic to human perception. This capability is crucial for numerous applications, ranging from autonomous driving [21, 36] and robotics [25, 26, 43] to augmented reality and content creation [58, 32, 2, 84], where an accurate understanding of dynamic scenes directly impacts system performance and user experience.
@@
Current neural-based video representations [44, 55, 81] predominantly rely on 2D and 2.5D techniques, treating videos as collections of spatiotemporal pixels [18, 89]. Although such discrete representations allow for basic temporal modeling through pixel matching [57] and tracking [28, 80], they struggle to capture complex scene dynamics and maintain appearance consistency, particularly in scenarios involving occlusions and non-rigid deformations [42, 44]. Moreover, they inherently lack explicit 3D understanding, leading to unreliable motion captures and spatial arrangements.
@@
Drawing inspiration from the fact that a monocular video is a projection of the dynamic 3D world, we ask: Can videos be represented in a canonical 3D space form without per-scene optimization? Recent advances in dynamic Gaussian Splatting [30] have shown remarkable capabilities in dynamic scene reconstruction [79, 74, 17] and video Gaussian representation [59, 70, 34], achieving high-fidelity rendering with explicit 3D representations, albeit through per-scene optimization. By modeling videos as flows of Gaussian primitives over time, we overcome the limitations of 2D representations and enable video Gaussian representation without per-scene optimization. As illustrated in Figure 1, this paradigm treats space-time holistically and offers key advantages: each Gaussian acts as a flexible building block that adapts to local appearance structures, enabling seamless representation of complex scenes. Moreover, when endowed with dynamic attributes, these structured Gaussians naturally approximate the underlying spatiotemporal volume of monocular videos, capturing intrinsic motions with temporal consistency and facilitating various downstream tasks such as object segmentation [53, 59], video editing [6, 46], and frame interpolation [15, 41].
@@
However, transforming casually captured monocular videos to dynamic Gaussian representations instantly (in seconds per frame) poses several challenges: (1) Unposed inputs. Gaussian Splatting [30] and its variants [85, 86, 24] heavily rely on accurate camera poses obtained through Structure-from-Motion (SfM) [64, 51], which are typically unavailable for casual monocular videos. Without such pose guidance, current methods [20, 4, 37, 72] fail to disentangle camera motion from scene dynamics, resulting in deteriorated rendering quality and even collapsed reconstruction. (2) Nonstructural Nature. Most existing dynamic Gaussian Splatting either leverage per-scene optimized deformation networks [79, 74, 17, 29, 33, 78] or adopt per-frame Gaussian generation [50] for dynamics modeling, both incompatible with our feed-forward prediction paradigm. The spatially unstructured property of Gaussian primitives further makes them prone to local minima in inverse rendering [94, 11], impeding feed-forward modeling of dynamic underlying structures in monocular videos. (3) Spatial Ambiguity. The absence of multi-view supervision and initialization from SfM points significantly limits the spatial modeling capability of Gaussian Splatting, leading to ambiguous scale, depth collapse, and inconsistent spatial arrangements in reconstructed scenes.
@@
To address these challenges, we introduce NutWorld to efficiently transform monocular videos into dynamic Gaussian Splatting representations in a single forward pass. Our method has three key components: (1) A structured spatial-temporal aligned Gaussian (STAG) representation in a canonical space (Section 4.1), enabling feed-forward prediction with pose-free, scale-invariant modeling. (2) An feed-forward pipeline (Section 4.2) that learns spatial-temporal correspondences and motions across frames, swiftly transforming them into STAG representations. (3) Depth and flow regularization (Section 4.3), leveraging calibrated depth [8, 77] and optical flow priors [75] to resolve spatial ambiguity and motion-appearance entanglement in the ill-posed monocular setting [59, 68]. With large-scale pre-training, NutWorld processes arbitrarily long videos while preserving spatial-temporal consistency through segment-based inference (Section 4.4).
@@
We performed qualitative and quantitative experiments on RealEstate10K [93] and MiraData [27] to verify the efficacy of NutWorld in video reconstruction. Moreover, our method demonstrates real-time inference speed and flexibility across various downstream tasks, including novel view synthesis, consistent depth estimation, video segmentation, video editing, and frame interpolation, indicating its potential as a general-purpose video representation framework. Our main contributions include:
@@
  • We present the first framework to efficiently represent world dynamics in casually captured monocular videos via dynamic Gaussian Splatting in a single forward pass.
@@
  • Extensive experiments on video reconstruction and various downstream tasks confirm the spatial-temporal coherence and versatility of NutWorld.
@@ 2 Related Work
Neural video representation. Casually captured monocular videos are widely available and can be considered as 2D projections of dynamic 3D scenes. Efficient representations of these videos are crucial for various computer vision tasks, including video object segmentation, object tracking, depth estimation, and frame interpolation. Early works [9, 55, 61, 44] leveraged implicit neural representations, modeling images through coordinate-based Multilayer Perceptrons (MLPs) through per-image optimization. Later works expanded upon this by incorporating learnable deformation field [44, 67, 40, 81], reconstructing videos as canonical and deformable MLPs. However, these methods often struggle to capture complex motions due to the absence of explicit 3D structure. Recently, explicit Gaussian video representations [68, 59, 70, 34] have emerged to address these limitations, representing videos explicitly as Gaussians in a 3D canonical space, each associated with a 3D motion projected onto the 2D video frames. But these methods still require computationally expensive per-video optimization, limiting their practical application. In contrast, our NutWorld introduces a novel feed-forward Gaussian video representation, distinguishing itself from previous approaches. The NutWorld network, trained on video datasets, efficiently reconstructs videos as Gaussians through a single forward pass, delivering improved reconstruction quality with substantial speedup.
Feed-Forward Gaussian Splatting. Recent advance in large-scale 3D scene datasets [38, 93, 39] has enabled feed-forward Gaussian approaches [7, 10, 88, 91, 73, 83, 52], which excel in efficiency and sparse view reconstruction. PixelSplat [7] and LatentSplat [73] employ the epipolar transformer [22, 71] to establish cross-view correspondences and predict 3D Gaussians from multi-view image features, while MVSplat [10] utilizes cost volumes to jointly predict depth and Gaussian parameters. Alternatively, GS-LRM [91] and Long-LRM [95] consider Gaussian Splatting reconstruction as a sequence-to-sequence translation task, employing transformer-based [65] or hybrid [13] architectures to regress Gaussian primitives. However, these feed-forward reconstruction methods designed for static 3D scenes have limitations when generalized to unconstrained videos, primarily because accurate per-frame camera poses cannot be obtained even through advanced pose estimation methods [66, 69, 90, 35]. To address this, we introduce a canonical camera space for building our NutWorld model, enabling robust 3D representation of dynamic scenes.
@@ 4.1 Spatial-Temporal Aligned Gaussian
Canonical camera space. Given an unposed monocular video, we employ an orthographic camera coordinate system to interpret the input as a quasi-3D canonical volume [67, 59] rather than an absolute 3D world space. This choice addresses two challenges: (1) the difficulty of obtaining consistent camera trajectories in dynamic scenes [64, 51, 45, 69, 66], and (2) the scale ambiguity in feed-forward 3D reconstruction [7, 10, 88], where perspective projection couples object size with camera distance. By imposing a fixed pose along the $z$ axis, orthographic projection removes perspective-induced distortions and eliminates the need for explicit camera estimation, enabling joint modeling of both camera and object motion. We detail the orthographic rasterization pipeline in the Appendix.
Structured Dynamic Gaussian. To overcome the unstructured nature in dynamic Gaussian Splatting and facilitate neural network integration, we introduce Spatial-Temporal Aligned Gaussian Splatting (STAG) within this canonical volume. STAG constrains each dynamic Gaussian to a specific pixel location and timestamp, in contrast to the previous approach of predicting unconstrained Gaussians with deformable fields across orthogonal space-time. Formally, for an input frame $F_{k}$ with normalized timestamp $t_{k}\in[0,1]$, we compute a Gaussian feature map $\mathcal{E}^{k}\in\mathbb{R}^{U\times V\times T}$, where $U$ and $V$ represent spatial dimensions, and $T$ denotes the channel dimension. Each $T$-dimensional pixel is decoded into a 3D Gaussian with an associated deformation field $\{\mathcal{G}_{i},\mu_{i}(t)\}$** Note that based on our empirical observation, we simplify the deformation modeling by considering only the center position, such that the deformation field reduces to $\mathcal{D}_{i}(t)=\mu_{i}(t)$. in a pixel-aligned manner [7, 60].
@@
As shown in Figure 2, the temporal slicing function $\mathbbm{1}(t_{k},t_{j})$ modulates the deformation field $\mu^{k}(t_{j})$ based on the temporal relationship between the rendering timestamp $t_{j}$ and the Gaussian’s reference timestamp $t_{k}$. For $t_{j}=t_{k}$, the deformation is suppressed ($\mathbbm{1}(t_{k},t_{j})=0$), preserving the original Gaussian position $\mu^{k}$ to maintain spatial alignment. When $t_{j}\neq t_{k}$, the deformation field is activated to adapt $\mu^{k}$ according to temporal differences, enabling per-pixel alignment across frames and enhancing consistency in our quasi-3D volume.
@@ 4.2 Encapsulate Dynamics within “Nutshell”
This mapping generates a set of STAGs $\{\mathcal{G}_{i},\mu_{i}(t)\}$ that can be rendered into arbitrary frames ($M\geq K$) through temporal interpolation of the deformation field. Specifically, the NutWorld model is composed of two main components:
Transformer-based Encoder. For each input frame $F_{k}\in\mathbb{R}^{H\times W\times 3}$, we augment pixels by concatenating their RGB values with corresponding depth coordinates $d^{*}_{k}$ along the channel dimension, obtained from a pre-trained video depth estimation model [8]. The augmented frames are first split into non-overlapping patches of size $p\times p$ using convolutions, which are then linearly transformed and concatenated across all frames to generate transformer input tokens. Notably, our architecture eliminates the need for explicit positional embeddings as used in ViT [16, 1], since depth coordinates inherently encode spatial information. The transformer blocks, comprising self-attention [65] and MLP layers, process these concatenated tokens to capture spatiotemporal correspondence and produce encoded features $\mathcal{E}_{0}\in\mathbb{R}^{K\times h\times w\times C}$, where $h=H/p$ and $w=W/p$ denote the spatial resolution and $C$ denotes the feature dimension. To ensure sufficient STAGs for casual videos, we leverage a hierarchical upsampling network [91, 76] that progressively expands the spatial resolution of the encoded feature $\mathcal{E}_{0}$. Each upsampler block first expands the feature dimension by a factor of 4 through a linear layer, followed by a PixelShuffle [54] layer that doubles the spatial resolution. The resulting features are then processed by a local attention layer with window size $\mathcal{W}$, which balances computational efficiency and spatial-temporal feature aggregation:
@@ 4.3 Calibrated 2D Priors Regularization
Learning spatially-aware STAGs solely from monocular videos is inherently ill-posed due to depth ambiguity and motion uncertainty. Therefore, we leverage off-the-shelf foundation models [75, 8] to recover spatial relationships and temporal motion through calibrated depth and flow priors, respectively. Note that in the quasi-3D canonical volume under the orthographic projection, the movement of STAG along the xy-coordinates directly corresponds to optical flow magnitude, whereas depth-related loss exclusively affects the z-coordinate, further facilitating the incorporation of those 2D priors.
Depth Regularization. To enhance robustness against scale and shift variations in depth rendering, we employ a scale and shift invariant loss [3, 49]. This loss function computes optimal scaling $\beta$ and shifting $\gamma$ factors that align the rendered depth $d$ with the pseudo depth $d^{*}$ estimated by the video depth prior [8]. The optimal values for $\beta$ and $\gamma$ are obtained by minimizing the squared error between the scaled rendered depth and the pseudo depth, as follows:
@@ 5.1 Experimental Setup
Training Dataset. NutWorld is pre-trained on MiraData [27] and RealEstate10K [93]. MiraData is a high-quality video dataset consisting primarily of 3D engine-generated scenes and movie clips with diverse motion patterns. The RealEstate10K dataset contains indoor house tour videos that showcase various architectural scenes and camera movements.  Unlike previous Generalizable 3DGS approaches [10, 7, 88], we treat RealEstate10K as a pure video dataset rather than a multi-view 3D scene dataset, thus not utilizing the COLMAP-calibrated camera poses. During pre-processing, we segment the original videos into video cubes, each containing 10 consecutive frames as the basic processing unit. Detailed information on the description of the dataset, preprocessing, and splitting of trains is provided in the Appendix.